Simplify: cos^2(34)- sin^2(34)
Please show the steps alongside it.

Answer by lwsshak3(11628)

cos^2(34)-sin^2(34)
Identity: cos(2x)=cos^2(x)-sin^2(x)
cos^2(34)-sin^2(34)=cos(2*34)=cos(68)